OpenStudy (anonymous):

The Marco Movie Theatre has 20 seats in the first row, 22 in the second, 24 in the third and so on for a total of 19 rows. How many total seats are in the Theater?

OpenStudy (anonymous):

You have the formula?

OpenStudy (anonymous):

tn=t1+(n-1)d Sn=n/2(t1+tn) formulas I got

OpenStudy (anonymous):

20=t1

OpenStudy (anonymous):

2= common difference

OpenStudy (anonymous):

I guess 19= N?

OpenStudy (anonymous):

2 I think could= D

OpenStudy (anonymous):

Formula I got

OpenStudy (anonymous):

do u have to use that formula, coz I got another way

OpenStudy (anonymous):

Nah you dont

OpenStudy (anonymous):

An=a1+(n-1)d

OpenStudy (anonymous):

20+(19-1)2

OpenStudy (anonymous):

=56

OpenStudy (anonymous):

now you use that other formula with 56

OpenStudy (anonymous):

S(19)=19(20+56)/2

OpenStudy (anonymous):

722 seats

thomaster (thomaster):

So why do you need help if you can do it yourself xD 20+22+24+26+28+30+32+34+36+38+40+42+44+46+48+50+52+54+56=722

OpenStudy (anonymous):

Cause I wonted a second opinion to see if I was right lol

thomaster (thomaster):

yea you're right :P

OpenStudy (anonymous):

Ill give you a medal for proving